OpenStudy (anonymous):

Let s and t be odd integers with x > t >= 1 and gcd(s,t) = 1. Prove that st, (s^2-t^2)/2, and (s^2+t^2)/2 are pairwise relatively prime.

ganeshie8 (ganeshie8):

have you tried anything yet

OpenStudy (anonymous):

I don't even know where to start...

OpenStudy (anonymous):

FYI, those three numbers are the formulas for generating primitive pythagorean triples, with a = st, b = (s^2-t^2)/2, and c = (s^2+t^2)/2. Also, a will be odd and b will be even.

ganeshie8 (ganeshie8):

\(\large \gcd(s, t) = 1 \iff\gcd(s^2, t^2) = 1\) agree ?

OpenStudy (anonymous):

Yes.

ganeshie8 (ganeshie8):

Next suppose \(\large d| st\) and \(\large d | \dfrac{s^2+t^2}{2}\) that means any prime factor\(\large p\) of \(d\) divides either \(\large s\) or \(\large t\), but not both because \(\large \gcd(s, t) = 1\) Suppose \(\large p | s \implies s = pj\), So \(\large p | \dfrac{(pj)^2+t^2}{2} \) but this is not possible because \(\large p \not | ~t\)

ganeshie8 (ganeshie8):

you can work the remaining pairs similarly

OpenStudy (anonymous):

I don't fully understand this, so I will have to study it.

ganeshie8 (ganeshie8):

sure :) you may ask me if there is a specific question

OpenStudy (anonymous):

Okay, thanks. If I ask a question after clicking best response, will you still get a notification, or does that close the problem out?

OpenStudy (anonymous):

If I click close and ask a question, will you still get a notification?

ganeshie8 (ganeshie8):

yes for all above questions ^^

OpenStudy (anonymous):

Cool, thanks for the help. I think my brain is tired tonight, so I will probably study this further tomorrow. :)

ganeshie8 (ganeshie8):

I get notification everytime you reply here/tag me. It won't depend on the status of question

ganeshie8 (ganeshie8):

np :) have good sleep !

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

I understand, thanks!